During testing, a formula 1 car of mass 750kg accelerates up a slope with inclination of 1 in 8. The car increases its velocity from (q) m/s to 48m/s and travels up the road a distance of (10g)m against a frictional resistance force of 1.0kN.


Using the Conservation of Energy principle, determine the following


parameters of the F1 racing car


a) The work done during the acceleration period


1
Expert's answer
2021-04-13T06:28:23-0400

"A=Fs=(ma-F_{fr})s=(\frac{mv^2}{2s}-F_{fr})s=\frac{mv^2}{2}-F_{fr}s,"

"A=\frac{750\cdot 48^2}{2}-1000\cdot10=854~kJ."